# Stellar Spectroscopy Using Diffraction Grating, CMOS Monochrome Sensor, and Reflecting Telescopes

### v1

@@ 1 Introduction
Since the 19th century, stellar spectroscopy has revolutionized our understanding of astrophysics, beginning with Fraunhofer’s discovery of dark absorption lines in the solar spectrum (Fraunhofer, 1814; Kirchhoff, 1860a, b). By the early 20th century, Annie Jump Cannon’s Harvard classification and Cecilia Payne’s work on stellar atmospheres established spectroscopy as the foundation of stellar astrophysics, linking spectral signatures to temperature, chemical composition, and stellar evolution (Payne-Gaposchkin, 1925; Cannon et al., 1897).
Today, high-resolution spectroscopy remains a cornerstone of observational astrophysics, enabling precise measurements of stellar abundances, radial velocities, and the detection of exoplanetary atmospheres (Gray, 2005; Martins et al., 2005). Recent usage of HIRES on the Keck I Telescope, employs a cross-dispersed echelle-grating spectrograph, achieving spectral resolutions between 25000 to 85000 across a wavelength range of 0.3 to 1.0 microns (Vogt et al., 1994). This high resolution allows for precise measurements of stellar parameters and has been instrumental in various astrophysical studies. Inspired by these sophisticated instruments, our project aims to construct a compact, low-cost diffraction-grating spectrometer integrated with a reflecting telescope and a CMOS sensor.
While operating at a lower resolution, the instrument embodies the similar fundamental techniques, providing a hands-on platform to observe and analyze stellar spectra across multiple spectral classes. This approach not only offers practical experience in astronomical instrumentation but also contributes to the growing field of accessible astronomical spectroscopy. Using a 600 lines/mm diffraction grating, a CMOS monochrome sensor, and a custom 3D-printed mount coupled to an 11-inch Schmidt-Cassegrain telescope, we constructed a spectrometer capable of dispersing starlight into well-defined spectra. Our targets – Sirius (A1V, the brightest star in the night sky), Vega (A0V, a standard photometric reference), Procyon (F5IV, a nearby subgiant), Capella (G8III + G0III, a quadruple system with a primary-pair of giants), and Betelgeuse (M2Ia, a red supergiant) (Gray et al., 2003; Keenan & McNeil, 1989; Kervella, P. et al., 2004; Torres et al., 2015) – were chosen to span a wide range of stellar types and evolutionary stages. Each of these stars carries historical significance in spectroscopy: Vega as the flux calibration standard, Sirius as an early benchmark in line-profile studies, and Betelgeuse as one of the first stars with a resolved stellar disk.
@@ 2.6 Data Processing Pipeline
Vega ($\alpha$ Lyrae, spectral type A0V) served as our primary spectrophotometric standard. We adopted the Hubble Space Telescope absolute spectrophotometry of Vega from Bohlin et al. (2014) as our reference spectrum. This choice addresses the lack of manufacturer-provided sensor response curves by implicitly calibrating instrumental sensitivity through comparison with a well-characterized stellar standard.
@@ 3.1 Spectral Features by Stellar Type
Capella (G8III + G0III) (Figure 7) exhibited further weakening of hydrogen features and strengthening of metallic lines. The Ca-II, H and K resonance lines appeared particularly strong, as expected for evolved giant stars with extended atmospheres. Multiple Fe-I and Fe-II lines dominated the blue-green spectral region (4000-5500 Å, Figure 7). The overall spectral energy distribution peaked in the yellow-green region, consistent with Capella’s composite surface temperature of $\sim$4970 K(Torres et al., 2015).
@@ Acknowledgements
The raw spectral data and the processed spectral datasets obtained using the telescope for stellar observations and calibration sources are available from the authors upon reasonable request. The data processing pipeline and analysis code used in this work are publicly available in our GitHub repository at https://github.com/niti21/P441.

### v2

@@ 1 Introduction
Since the 19th century, stellar spectroscopy has revolutionized our understanding of astrophysics, beginning with Fraunhofer’s discovery of dark absorption lines in the solar spectrum (fraunhofer1814; Kirchoff1860FL; Kirchoff1860Sun). By the early 20th century, Annie Jump Cannon’s Harvard classification and Cecilia Payne’s work on stellar atmospheres established spectroscopy as the foundation of stellar astrophysics, linking spectral signatures to temperature, chemical composition, and stellar evolution (1925Payne; maury1897spectra).
Today, high-resolution spectroscopy remains a cornerstone of observational astrophysics, enabling precise measurements of stellar abundances, radial velocities, and the detection of exoplanetary atmospheres (2005DavidG; 2005Martin). Recent usage of HIRES on the Keck I Telescope, employs a cross-dispersed echelle-grating spectrograph, achieving spectral resolutions between 25000 to 85000 across a wavelength range of 0.3 to 1.0 microns (HIRES). This high resolution allows for precise measurements of stellar parameters and has been instrumental in various astrophysical studies. Inspired by these sophisticated instruments, our project aims to construct a compact, low-cost diffraction-grating spectrometer integrated with a reflecting telescope and a CMOS sensor.
While operating at a lower resolution, the instrument embodies the similar fundamental techniques, providing a hands-on platform to observe and analyze stellar spectra across multiple spectral classes. This approach not only offers practical experience in astronomical instrumentation but also contributes to the growing field of accessible astronomical spectroscopy. Using a 600 lines/mm diffraction grating with a CMOS monochrome sensor, and a custom 3D-printed mount coupled to an 11-inch Schmidt-Cassegrain telescope (Figure 1), we constructed a spectrometer capable of dispersing starlight into well-defined spectra. Our targets – Sirius (A1V, the brightest star in the night sky), Vega (A0V, a standard photometric reference), Procyon (F5IV, a nearby subgiant), Capella (G8III + G0III, a quadruple system with a primary-pair of giants), and Betelgeuse (M2Ia, a red supergiant) (Gray_2003_Sir_Vega; Perkins_BG_cap; procyon0; Torres_2015_Capella_bin) – were chosen to span a wide range of stellar types and evolutionary stages. Each of these stars carries historical significance in spectroscopy: Vega as the flux calibration standard, Sirius as an early benchmark in line-profile studies, and Betelgeuse as one of the first stars with a resolved stellar disk.
@@ 2.6 Data Processing Pipeline
Vega ($\alpha$ Lyrae, spectral type A0V) served as our primary spectrophotometric standard. We adopted the Hubble Space Telescope absolute spectrophotometry of Vega from bohlin2014 as our reference spectrum. This choice addresses the lack of manufacturer-provided sensor response curves by implicitly calibrating instrumental sensitivity through comparison with a well-characterized stellar standard.
@@ 3.1 Spectral Features by Stellar Type
Capella (G8III + G0III) (Figure 7) exhibited further weakening of hydrogen features and strengthening of metallic lines. The Ca-II, H and K resonance lines appeared particularly strong, as expected for evolved giant stars with extended atmospheres. Multiple Fe-I and Fe-II lines dominated the blue-green spectral region (4000-5500 Å, Figure 7). The overall spectral energy distribution peaked in the yellow-green region, consistent with Capella’s composite surface temperature of $\sim$4970 K(Torres_2015_Capella_bin).
@@ Acknowledgements
The raw spectral data and the processed spectral datasets obtained using the telescope for stellar observations and calibration sources are available from the authors upon reasonable request. The data processing pipeline and analysis code used in this work are publicly available in our GitHub repository at https://github.com/niti21/Stellar-Spectrometry.